# A Copula Based Supervised Filter for Feature Selection in Diabetes Risk Prediction Using Machine Learning

### v2

@@ Results
The Gumbel-$\lambda_{U}$ feature set shows a clear improvement over the other filter methods. Its ROC-AUC of 0.823 is higher than the scores for Mutual Information (MI) at 0.821 and mRMR at 0.794, which did not identify the most effective feature combinations for this dataset. This result indicates that focusing on upper-tail dependence helps our method find a more relevant predictive signal than traditional information-theoretic or nearest-neighbor methods.
@@
In contrast, McNemar’s test, applied to the paired correctness vectors, returned a highly significant p-value ($p<10^{-3}$) when comparing Gumbel vs. All, indicating the two models make different types of errors on the same instances, despite similar AUCs.
Table 4: Paired statistical tests on the CDC test set comparing the Gumbel-selected feature set against alternatives using the Gradient Boosting classifier. DeLong compares ROC–AUC; McNemar compares error profiles.
Comparison DeLong p-value McNemar $\chi^{2}$ McNemar p-value
Gumbel vs. All Features 0.999562 94.8574 $2.05\times 10^{-22}$
Gumbel vs. MI 0.999808 135.9645 $2.03\times 10^{-31}$
Gumbel vs. mRMR 0.996744 496.0476 $6.89\times 10^{-110}$
Gumbel vs. ReliefF 0.999909 70.2984 $5.10\times 10^{-17}$
Gumbel vs. L1/EN 0.999665 158.0189 $3.07\times 10^{-36}$
@@
To understand the relative contribution of each predictor within our top-performing feature set, we conducted a permutation importance analysis. We used the best-performing configuration, the Gradient Boosting (GB) model trained on the 10 Gumbel-selected features, and measured the drop in ROC–AUC on the held-out test set when each feature was randomly permuted.
Figure 3: Permutation importance (mean $\Delta$ ROC–AUC over 500 permutations) for the Gradient Boosting model trained on the 10 Gumbel-selected features (CDC test set). Larger values indicate greater importance.
@@
We stress–tested the Gradient Boosting (GB) classifier trained on the Gumbel top-10 features under three perturbations (Table 6). Relative to the baseline ROC–AUC of 0.823, performance remained within $\approx 0.001$ under label–flip 5% (0.8223) and feature noise $\sigma{=}0.10$ (0.8227), and declined by $\approx 0.011$ under MCAR 10% (0.8127), confirming robustness under light noise.
@@
Next, we present the classification results.
@@
The predictive performance for each feature ranking on the PIMA dataset is presented in Table 8. A unique characteristic of this benchmark is that it contains only eight predictors, so every method ultimately ranks the full feature set. The observed performance differences therefore arise from how each method’s ranking of those eight features interacts with the learning algorithm of the classifier.
@@
On this dataset, our Gumbel-$\lambda_{U}$ method achieved the highest overall performance. The top ROC-AUC of 0.867 was obtained by pairing the Gumbel-ranked features with a Random Forest (RF) classifier. This result slightly surpassed other strong combinations, including those from ReliefF (0.866 with GB) and L1EN (0.864 with GB). This outcome is a strong validation, demonstrating that even in a low-dimensional clinical setting where feature removal is not the primary goal, our method’s focus on upper-tail dependence identifies a feature ranking that is optimal for a powerful non-linear classifier like Random Forest.
To visualize the PIMA results, Figure 4 plots the superimposed ROC curves for the Random Forest model across all feature sets. The curves are tightly clustered, indicating very similar discrimination. The Gumbel-$\lambda_{U}$ set attains the highest AUC (0.867), edging the next best lines viz. mRMR, All features, L1EN, and ReliefF. MI is slightly lower. The small visual separations and the stair-step shape reflect the modest test size, but the plot confirms that our tail-dependence ranking is at least as strong as the baselines and, for RF, slightly superior.
@@
Takeaway:
In settings with a small number of features, where feature selection is not needed to manage model complexity, our upper-tail dependence ranking still performs at a high level. It highlights clinically relevant factors such as Glucose, BMI, and Age, supporting CDC findings that focusing on joint extremes can be informative.
To determine if the small differences between our method and the top baselines are statistically significant, we now present the results of paired statistical tests
@@
Next, we used DeLong’s test to compare the ROC-AUC of the best-performing classifier for our method (Random Forest) against the other feature sets. The results are reported in Table 9. The test yielded consistently high p-values (p ¿ 0.99 for all comparisons), leading us to fail to reject the null hypothesis. We conclude there is no statistically significant difference in the AUC scores. This is a crucial validation, as it confirms that our Gumbel-$\lambda_{U}$ ranking achieves a state-of-the-art performance that is statistically indistinguishable from the strongest baselines.
@@
Gumbel vs. All Features ¿ 0.998226 0.13 0.724
Gumbel vs. MI ¿ 0.9974 3.37 0.066
Gumbel vs. mRMR ¿ 0.998799 0.00 1.000
Gumbel vs. L1EN ¿ 0.997941 2.08 0.149
Gumbel vs. ReliefF ¿ 0.997656 2.78 0.095
Next, we look at the permutation importance within the Gumbel set.
@@
As above, the selector is $O(d\,n\log n)$; on the small PIMA split, all filter methods are within the same order of magnitude. We recorded the wall-clock time for each selector on the PIMA training split (Table 12). On this smaller dataset, L1/Elastic-Net is the fastest at 0.0049 s, while ReliefF remains the slowest at 0.4711 s. Crucially, our Gumbel-$\lambda_{U}$ method is comparably fast to other standard filters like MI and mRMR (all in the $\sim$0.02–0.03 s range), demonstrating that its state-of-the-art predictive performance on this dataset comes without any significant computational penalty.
@@
We tested our method on two different diabetes datasets and found strong support for its usefulness. On the large CDC public health dataset, our Gumbel-$\lambda_{U}$ selector was competitive and was the fastest method, running much quicker than other approaches. It also reduced the number of features by about $\approx$ 52% (from 21 to 10) without losing predictive accuracy compared to using all features. On the well-known PIMA clinical dataset, our method produced the best-performing model with the highest ROC-AUC among all tested options. Although it was not the fastest on this smaller dataset, its speed was still close to other standard filters, showing that its strong performance did not come at a high computational cost. In both cases, our method consistently found clinically important predictors, showing it can identify meaningful and understandable signals.

### v4

@@ Results
The Gumbel-$\lambda_{U}$ feature set shows a clear advantage over several other filter methods. Its ROC-AUC of 0.823 is numerically higher than the scores for Mutual Information (MI) at 0.821 and mRMR at 0.794. As the formal statistical tests in the next section will confirm, our method is not only competitive but statistically superior to these common filters, indicating that its focus on upper-tail dependence captures a more relevant predictive signal.
@@
Now we use a covariance-aware DeLong’s test to compare ROC-AUCs in a paired manner. The model trained on all 21 features achieved the highest performance (ROC-AUC = 0.827). As expected given the modest full set (21 variables) and our emphasis on parsimony, we focus comparisons among selectors that return exactly 10 features (k = 10).
@@
Finally, McNemar’s test on paired error profiles is highly significant for Gumbel vs. each baseline (Table 4), indicating the models make different types of mistakes even when AUCs are similar.
Note. Extremely small p-values are reported as $p<0.001$ instead of 0.
Table 4: Paired tests on the CDC test set comparing the Gumbel-selected feature set against alternatives (Gradient Boosting). DeLong compares ROC-AUC; McNemar compares error profiles.
Comparison DeLong $p$ McNemar $\chi^{2}$ McNemar $p$
Gumbel vs. All <0.001<0.001 94.86 $2.05\times 10^{-22}$
Gumbel vs. MI $1.47\times 10^{-5}$ 135.96 $2.03\times 10^{-31}$
Gumbel vs. mRMR <0.001<0.001 496.05 $6.89\times 10^{-110}$
Gumbel vs. ReliefF $0.154$ 70.30 $5.10\times 10^{-17}$
Gumbel vs. L1EN $2.69\times 10^{-9}$ 158.02 $3.07\times 10^{-36}$
@@
To understand which variables our method uses, we computed permutation importance for our best CDC setup: Gradient Boosting trained on the Gumbel top-10 features (Figure 3). On the held-out CDC test set, we permuted one feature at a time (500 repetitions) and measured the mean drop in ROC-AUC. Larger drops mean the model relies more on that feature.
[…]
@@
After this, we evaluated the stability of our Gumbel-selected feature set under several common data perturbations.
@@
We stress–tested the Gradient Boosting (GB) classifier trained on the Gumbel top-10 features under three perturbations (Table 6). Relative to the baseline ROC-AUC of 0.823, performance remained within $\approx 0.001$ under label–flip 5% (0.8223) and feature noise $\sigma{=}0.10$ (0.8227), and declined by $\approx 0.011$ under MCAR 10% (0.8127), confirming robustness under light noise.
@@
PIMA serves as a complementary clinical benchmark where no dimensionality reduction is possible (eight total predictors). Consequently, all methods use the same variables; differences reflect sampling and calibration randomness (column order has no effect for LR/RF/GB/XGB). Under this setup, the Gumbel ranking paired with Random Forest achieved the numerically highest ROC-AUC (0.867). Formal paired tests are reported in the Statistical Tests subsection.
@@
After this, we present the classification results.
@@
The predictive performance for each feature ranking on the PIMA dataset is presented in Table 8. A unique characteristic of this benchmark is that it contains only eight predictors, so every method ultimately ranks the full feature set. Because every model trains on exactly the same eight variables, AUCs are expected to be very close. The small differences observed reflect sampling/calibration randomness rather than meaningful effects, which, as later confirmed by our DeLong tests (all p ¿ 0.05), are not significant.
@@
On this dataset, our Gumbel-$\lambda_{U}$ ranking yielded the numerically highest ROC-AUC. The top value of 0.867 was obtained by pairing the Gumbel-ranked features with a Random Forest (RF) classifier. This edges strong alternatives (e.g., ReliefF: 0.866 with GB; L1EN: 0.864 with GB). In a low-dimensional clinical setting where feature removal is not the primary goal, this indicates that a tail-aware ranking is competitive while remaining interpretable, rather than conclusively superior.
To visualize the PIMA results, Figure 4 shows superimposed ROC curves for the Random Forest model across all feature sets. The curves are tightly clustered, indicating very similar discrimination. The Gumbel-$\lambda_{U}$ set attains the numerically highest AUC (0.867), with mRMR, All, L1EN, and ReliefF close behind and MI slightly lower.
@@
A key finding is that even with a small number of features where feature selection is not needed to manage model complexity, our upper-tail dependence ranking still performs at a high level. It highlights clinically relevant factors such as Glucose, BMI, and Age, supporting CDC findings that focusing on joint extremes can be informative.
Note: PIMA has 8 predictors; all methods use the same variables. Reported differences therefore arise from ranked ordering and model refitting/calibration rather than feature removal.
@@
Next, we used a covariance-aware DeLong’s test to compare the ROC-AUC of the best-performing classifier (Random Forest) trained on the Gumbel-$\lambda_{U}$ features against each baseline. As shown in Table 9, all pairwise comparisons yielded $p>0.05$. Therefore, we fail to reject the null in every case and conclude there is no statistically significant difference in AUC. This supports that, in this low-dimensional clinical setting, the Gumbel-$\lambda_{U}$ ranking achieves performance statistically indistinguishable from strong baselines.
@@
Gumbel vs. All Features 0.253 0.13 0.724
Gumbel vs. MI 0.095 3.37 0.066
Gumbel vs. mRMR 0.504 0.00 1.000
Gumbel vs. ReliefF 0.101 2.78 0.095
Gumbel vs. L1EN 0.255 2.08 0.149
After this, we look at the permutation importance within the Gumbel set.
@@
As above, the selector is $O(d\,n\log n)$; on the small PIMA split, all filter methods are within the same order of magnitude. We recorded the wall-clock time for each selector on the PIMA training split (Table 12). On this smaller dataset, L1/Elastic-Net is the fastest at 0.0049 s, while ReliefF remains the slowest at 0.4711 s. Crucially, our Gumbel-$\lambda_{U}$ method is comparably fast to other standard filters like MI and mRMR (all in the $\sim$0.02–0.03 s range), demonstrating competitive predictive performance on this dataset without any significant computational penalty.
@@
Limitation. PIMA’s low dimensionality limits conclusions about feature reduction; future work will assess richer multi-center EHR cohorts to evaluate tail-dependent selection under higher dimensionality and domain shift.
Two regimes, complementary evidence.
On CDC (21 features), upper-tail filtering reduces the feature set by $\sim$52%. This introduces a minor but statistically significant performance trade-off against using the full feature set, but our method’s key advantage is highlighted by its statistically superior performance compared to other filter baselines, demonstrating its efficiency and ranking power. On PIMA (8 features), where selection is not applicable, our criterion yields clinically coherent rankings and maintains discrimination (no significant AUC differences), i.e., it does no harm in low-dimensional clinical settings. Together, these results support the method’s robustness, interpretability, and practicality across dataset types.
@@
We tested our method on two different diabetes datasets and found strong support for its usefulness. On the large CDC public health dataset, our Gumbel-$\lambda_{U}$ selector was competitive and was the fastest method, running much quicker than other approaches. It also reduced the number of features by about $\approx$ 52% (from 21 to 10), and on this parsimonious set, it proved its value as a top-tier filter by delivering performance that was statistically superior to standard filters like Mutual Information and mRMR and indistinguishable from the strong ReliefF baseline. On the PIMA dataset (8 predictors), our ranking achieved the numerically highest ROC-AUC while remaining statistically indistinguishable from strong baselines, functioning as a ranking-only clinical sanity check. Although it was not the fastest on this smaller dataset, its speed was still close to other standard filters, showing that its strong performance did not come at a high computational cost. In both cases, our method consistently found clinically important predictors, showing it can identify meaningful and understandable signals.